Equipment was purchased for $87000 on January 1, 2021. Freight charges amounted to $2400 and there was a cost of $10000 for building a foundation and installing the equipment. It is estimated that the equipment will have a $16000 salvage value at the end of its 5-year useful life. What is the amount of accumulated depreciation at December 31, 2022 if the straight-line method of depreciation is used?

$33360.

$14920.

$28800.

$16680.
Answer #1

On Jan. 1, 2021

Cost of Equipment      = $87,000

Freight charges            = $2,400

Installation Cost          = $10,000

Salvage value              = $16,000

Useful life                   = 5 Years

Calculation of annual depreciation under straight line method of depreciation: -

= (87,000 + 2,400 + 10,000 – 16,000)/ 5

= 83,400/ 5

= $16,680

Accumulated Depreciation at Dec. 31, 2022: -

= 16,680*2

= $33,360

Hence, first option is correct.